# A multi-stage Bayesian approach to fit spatial point process models

### v1

@@ 3.1 Recursive Bayes Algorithm
We first show that the complete likelihood in (1) can be decomposed into two terms:
@@
Moreover, note that the first term in (7) is equivalent to the conditional likelihood in (2). The second term, the likelihood for $n$, follows a Poisson distribution:
@@
Evaluating the decomposed posterior distribution in (9) using PPRB results in two stages. In the first stage, we evaluate the transient posterior using the conditional likelihood (2) and the prior for $\boldsymbol{\beta}$
@@
to obtain a Bayesian sample, $\big{\{}\boldsymbol{\beta}^{(k)}|\{\mathbf{s}_{i}\},n\big{\}}_{k=1}^{K}$. Recall that in section 2.1.1, we show that fitting the conditional likelihood in (2) can be approximated using logistic regression, which permits a variety of ways to obtain a sample from the transient posterior distribution. We describe and compare first-stage sampling strategies in the context of the multi-stage algorithm in the following section.
In between stages, we use the transient posterior realizations to compute corresponding values of $\Lambda(\mathcal{S})$ in parallel and store them for recall in the following stage. Approximating $\Lambda(\mathcal{S})$ is the most computationally intensive step in evaluating the likelihood for $n$ and parallelization results in significant speedup for the final stage of the algorithm.
In the second stage, we use the transient posterior distribution as a prior and proposal distribution in an MCMC algorithm to account for the stochasticity associated with $n$ in the full posterior distribution. Using the transient posterior distribution (10) as a prior and proposal distribution for $\boldsymbol{\beta}$ results in convenient cancellations involving the data model conditioned on $n$ in the M-H acceptance ratio:
@@
Recall that we are unable to learn about $\beta_{0}$ using the first-stage logistic regression approximation. It is, however, crucial to efficiently propose $\beta_{0}$ for the M-H correction in the second stage to obtain reasonable MCMC mixing rates. Conveniently, if we transform the intercept such that $\zeta=e^{\beta_{0}}$, we can obtain a Gibbs sampler using a $\text{Gamma}(a,b)$ prior for $\zeta$. The Gibbs update for $\zeta$ becomes
@@
We then log-transform our posterior sample for $\zeta$ to obtain a sample from the $\beta_{0}$ full-conditional distribution. For a full derivation of the Gibbs update for $\zeta$, see Appendix A.
The Gibbs sampler for $\beta_{0}$ also allows for the M-H ratio in (13) to simplify to
@@
Thus, each MCMC iteration in the second stage only requires updating $\beta_{0}$ using the Gibbs update in (14) and evaluating $\big{[}n|{\beta_{0}}^{(k)},\boldsymbol{\beta}^{(*)}\big{]}$, both of which are fast using the pre-computed values of $\Lambda(\mathcal{S})$ from the intermediate stage. The full multi-stage algorithm is summarized in Algorithm 1.
@@
Algorithm 1 Fully Bayesian Multi-stage MCMC
@@ 3.2 First-stage Sampling Strategies
One advantage to fitting the Bayesian SPP model in multiple stages is the increased flexibility facilitated by approximating the first stage using a logistic regression model. There are a variety of approaches for fitting a Bayesian logistic regression model such as the Pólya-Gamma data augmentation method described by Polson et al. (2013), which results in a fully Gibbs algorithm and circumvents the need to perform parameter tuning. Other unsupervised Monte Carlo methods for fitting logistic regression can be performed using existing software, such as stan_glm for Hamiltonian Monte Carlo (HMC; Neal, 2011). We refer to fitting the multi-stage procedure with the Pólya-Gamma data augmentation and stan_glm approaches as the Pólya-Gamma (PG) and Bayesian GLM (B-GLM) methods, respectively, in reference to their first-stage sampling strategies.
An alternative to fitting the multi-stage algorithm as fully Bayesian is to fit the first stage using maximum likelihood estimation. This can be attractive because fitting a Bayesian logistic regression model in the first stage can be slow, especially after performing large data augmentations, such as the background points generated from the Berman-Turner device and/or the latent variables required for the Pólya-Gamma Gibbs sampler. This creates a computational bottleneck because the first stage cannot be parallelized and thus does not scale with the number of computational cores. Johnson et al. (2022) proposed a multi-stage algorithm for a different class of models in which the transient posterior distribution is approximated using non-Bayesian likelihood maximization in the first stage. A sample from the approximated transient posterior distribution can then be corrected using an approximate M-H ratio in later stages (Johnson et al., 2022).
Applying a similar approach, we can perform fast maximum likelihood estimation for $\boldsymbol{\beta}$ in the first stage using the glm function in R to fit the logistic regression model described in section 2.2.1. If we assume a flat prior for $\boldsymbol{\beta}$ and standard regularity conditions (i.e., the likelihood is continuously twice-differentiable), we can apply large sample theory to obtain the following approximation:
@@
Johnson et al. (2022) showed that the approximate transient posterior distribution in (16) allows for approximate cancellations in the later stage M-H ratio and that parameter estimates obtained using this method are close to those obtained through a conventional single-stage MCMC algorithm. This method can be applied in our multi-stage procedure to obtain an approximate and fast sample from the posterior distribution for $\boldsymbol{\beta}$. Using the Gibbs sampler in (14), the M-H ratio in the second stage again reduces to (15). We refer to this multi-stage method as the GLM approximate (GLM-A) method hereafter.
If more exact estimates are desired, the second-stage M-H correction can be amended to account for the transient posterior approximation in (16). Instead of decomposing the complete likelihood and making approximate cancellations like in (13), we can use the original M-H correction as in (11). Using the Gibbs sampler for $\zeta$ based on the full-conditional distribution in (14) results in the following M-H ratio:
@@
To speed up the evaluation of the log-likelihood for $\big{[}\{\mathbf{s}_{i}\},n|{\beta_{0}}^{(*)},\boldsymbol{\beta}^{(*)}\big{]}$, we compute $\sum_{i=1}^{n}\mathbf{x}^{\prime}(\mathbf{s}_{i})\boldsymbol{\beta}$ in parallel in the second stage in addition to $\Lambda(\mathcal{S})$ using the first stage sample. Although the intermediate and second stages involve more calculations compared to the previous multi-stage methods, the overall algorithm results in significant speedup for this model compared to the conventional single-stage method because of the fast first-stage evaluation using maximum likelihood estimation and the ability to parallelize large computations over $k$ in the intermediate stage. We refer to this multi-stage method as the GLM exact (GLM-E) method hereafter.
Both the GLM-A and GLM-E methods result in significant speedup when compared to both the single-stage MCMC algorithm and fully Bayesian multi-stage methods (e.g., PG and B-GLM). Furthermore, the independent samples from the first stage can help prevent sample degeneracy and improve MCMC mixing in the second stage. Table 1 summarizes the strengths of each first-stage sampling strategy.
Property PG B-GLM GLM-E GLM-A
@@
Exact M-H $\times$
Independent first stage samples $\times$ $\times$
@@ 3.3.2 Posterior Point Simulation
We can simulate posterior predictive event locations in the unobserved region $\mathcal{S}_{0}$ after obtaining estimates for $\lambda(\mathbf{u})$ $\forall\mathbf{u}:\mathbf{u}\in\mathcal{S}_{0}$ using the Lewis-Schedler method for simulating from an IPP (Lewis and Schedler 1979). Alternatively, a new point realization for the entire study domain $\mathcal{S}$ can be simulated using posterior estimates for $\lambda(\mathbf{u})$ $\forall\mathbf{u}:\mathbf{u}\in\mathcal{S}$. These results can then be used to characterize posterior predictive derived quantities.
@@ 4 Simulation Study
We demonstrate the multi-stage MCMC approach using a simulation study in a compact window setting. The data were first generated from an IPP with an intensity function that is log-linear in terms of three known parameters ($\beta_{0}=5$, $\beta_{1}$ = 1, $\beta_{2}$ = 2) across the entire study domain (Figure 1a). This resulted in the generation of $N=825$ total events of which $n=474$ were located within the compact observation windows (Figure 1b).
Figure 1: (a) Intensity heat map for simulated data. (b) Realization of simulated points. Points located within the compact windows (red; $n$ = 474) were used to fit the model.
Only the events located within the compact observation windows were used to fit the model with the windowed complete likelihood (3). We fit the simulated data using the first-stage sampling strategies described in section 3.2 and a conventional single-stage MCMC algorithm for comparison on a 24-core machine with 3.68 GHz processors and 192 GB of RAM. Figure 2 compares the posterior samples for $\beta_{0}$ and $\boldsymbol{\beta}$ obtained using the single-stage and the multi-stage MCMC algorithms. It is evident that the marginal posterior distributions for $\beta_{0}$ and $\boldsymbol{\beta}$ are very similar among implementation approaches. Table 2 compares the seconds per effective sample size between different methods for each parameter. The time recorded for the single-stage method does not include a priori parameter tuning, while the remaining multi-stage methods did not require tuning. Both the non-Bayesian first-stage methods showed an order of magnitude improvement. This is due to an improvement in both runtime and MCMC effective sample size. Figure 3 shows the posterior predictive distribution for $N$ obtained using the posterior samples from the GLM-E multi-stage method.
@@ 5 Application: Harbor Seal Pups in Johns Hopkins Inlet
Previous studies analyzing these data have focused on modeling harbor seal abundance for the entire inlet or using coarse spatial grids. For example, Womble et al. (2020) incorporated the aerial imagery survey data along with other sources of data to obtain estimates and uncertainty quantification for harbor seal abundance. Womble et al. (2021) fit a multivariate conditional autoregressive (MCAR) model to summarize spatially explicit abundance on a 200 m $\times$ 300 m grid ($L=434$ total grid cells). Jansen et al. (2015) used harbor seal data from a similar aerial survey collected in 2002 to fit a generalized additive model (GAM) and summarized spatially explicit abundance on a grid of 400 m $\times$ 400 m cells ($L=779$ total grid cells). Previous studies on spatially explicit abundance of harbor seals in Johns Hopkins Inlet have not explored modeling the locations as a spatial point process directly. Modeling these data as an SPP provides unique advantages because it allows for informed simulation of potential seal locations outside the image footprints using an established statistical framework. Furthermore, the multi-stage algorithm makes fitting the model on a fine spatial grid more computationally flexible and feasible, especially with the availability of distributed computing resources.
We fit a Bayesian SPP model to harbor seal pup haul out location data from 21 June 2007 ($n=452$; Figure 5) using the aforementioned multi-stage MCMC approaches. We modeled harbor seal pups because they are the most vulnerable segment of the population and are especially susceptible to disturbance from vessels (Mathews et al., 2016). Furthermore, pups are hauled out on ice and observable with high probability at the time the survey was conducted (i.e., in the month of June and in the early afternoon) based on secondary analysis of wet-dry tag data of harbor seal pups in Disenchantment Bay, Alaska. Our study domain spanned from the Johns Hopkins Glacier terminus to Jaw Point, which is approximately 9km long and 1.5-2km wide (Figure 4a). The image footprints for this survey cover approximately half of the study domain. We assume perfect detection of hauled out pups within the image footprints and no detection outside the image footprints, giving rise to compact observation windows. Additionally, due to the systematic surveying within the study domain, we assume the unobserved areas in the inlet to be ignorable missing data (i.e., the covariate values and relationship between covariates and pup locations do not differ significantly between observed and unobserved areas). Using the pup locations as point events, we fit the SPP model using the windowed complete likelihood in (3).
@@
We fit the model using three covariates: ice proportion, bathymetry, and distance to glacier terminus (Appendix B). The ice data were represented as iceberg outlines within the image footprints and summarized as ice proportions to allow for spatial prediction outside the image footprints (Appendix C). Each covariate was summarized on the same raster grid with a spatial resolution of approximately 10 m $\times$ 10 m ($L=304,945$ total grid cells). The grid centers were used as quadrature points throughout the analysis.
5.1 Neural Network Basis Expansion
To increase model flexibility and allow for more complex relationships between covariates and intensity, we constructed a basis representation of $\mathbf{X}$, the $p$ $\times$ $L$ design matrix consisting of the covariate values at each grid center. Of the basis functions typically used to account for spatial heterogeneity (e.g., fixed-rank basis functions and radial basis functions; Hefley et al., 2017; Wikle and Zammit-Mangion, 2023), we found that a single-hidden layer feedforward neural network (SLFNN) was reasonably able to capture the relationship between the covariates and pup space use. Additionally, we found that the Extreme Learning Machine (ELM) scheme (Huang et al., 2006) was sufficient for constructing the neural network (NN) weights and resulted in efficient computation in the multi-stage algorithm.

### v2

@@ 3.1 Recursive Bayes Algorithm
We first show that the complete likelihood (1) can be decomposed into two terms:
@@
Moreover, note that the first term in (7) is equivalent to the conditional likelihood (2). The second term, the likelihood for $n$, follows a Poisson distribution:
@@
Evaluating the decomposed posterior distribution (9) using PPRB results in two stages. In the first stage, we evaluate the transient posterior using the conditional likelihood (2) and the prior for $\boldsymbol{\beta}$
@@
to obtain an MCMC sample, $\big\{\boldsymbol{\beta}^{(k)}|\{\mathbf{s}_{i}\},n\big\}_{k=1}^{K}$. Recall that in section 2.2.1, we show that fitting the model using the conditional likelihood (2) can be approximated using logistic regression, which permits a variety of ways to obtain a sample from the transient posterior distribution. We describe and compare first-stage sampling strategies in the context of the multi-stage algorithm in the following section.
In the intermediate stage, we use the transient posterior realizations to compute corresponding values of $\Lambda(\mathcal{S})$ in parallel and store them for recall in the following stage. Approximating $\Lambda(\mathcal{S})$ is the most computationally intensive step in evaluating the likelihood for $n$ and parallelization results in significant speedup for the final stage of the algorithm.
In the second and final stage, we use the transient posterior distribution as a prior and proposal distribution in an MCMC algorithm to account for the stochasticity associated with $n$ in the full posterior distribution (4). Using the transient posterior distribution (10) as a prior and proposal distribution for $\boldsymbol{\beta}$ results in convenient cancellations involving the data model conditioned on $n$ in the M-H acceptance ratio:
@@
Recall that we are unable to learn about $\beta_{0}$ using the first-stage logistic regression approximation. It is, however, crucial to efficiently propose $\beta_{0}$ for the M-H correction in the second stage to obtain reasonable MCMC mixing rates and prevent sample degeneracy. Conveniently, if we transform the intercept such that $\zeta=e^{\beta_{0}}$, we can obtain a Gibbs sampler using a $\text{Gamma}(a,b)$ prior for $\zeta$. The Gibbs update for $\zeta$ is based on the full-conditional distribution
@@
We then log-transform our posterior sample for $\zeta$ to obtain a sample from the full-conditional distribution $\beta_{0}|\cdot]$. For a full derivation of the Gibbs update for $\zeta$, see Appendix A.
The Gibbs sampler for $\beta_{0}$ also allows for the M-H ratio (13) to simplify to
@@
Thus, each MCMC iteration in the second stage only requires updating $\beta_{0}$ using the Gibbs update (14) and evaluating the Poisson likelihood $\big[n|{\beta_{0}}^{(k)},\boldsymbol{\beta}^{(*)}\big]$, both of which are fast using the pre-computed values of $\Lambda(\mathcal{S})$ from the intermediate stage. The resulting MCMC realizations from the second stage are realizations from the full posterior distribution (4). The full multi-stage algorithm is summarized in Algorithm 1.
@@
Algorithm 1 Multi-stage PPRB MCMC
@@ 3.2 First-stage Sampling Strategies
One advantage to fitting the Bayesian SPP model in multiple stages is the increased flexibility facilitated by the ability to approximate the first stage using a logistic regression model. There are a variety of approaches for fitting a Bayesian logistic regression model such as the Pólya-Gamma data augmentation method described by Polson et al. (2013), which results in a fully Gibbs algorithm and alleviates the need to perform parameter tuning. Other unsupervised Monte Carlo methods, such as HMC with NUTS, for fitting logistic regression can be performed using existing software, such as rstanarm (Gabry and Goodrich, 2016) or brms (Bürkner et al., 2025) in R. We refer to fitting the multi-stage procedure with the Pólya-Gamma data augmentation and HMC approaches as the Pólya-Gamma (PG) and HMC methods, respectively, in reference to their first-stage sampling strategies.
An alternative to fitting the multi-stage algorithm as fully Bayesian is to fit the first stage using maximum likelihood estimation. This can be attractive because fitting a Bayesian logistic regression model in the first stage can be slow, especially after performing large data augmentations, such as the background points generated from the Berman-Turner device and/or the latent variables required for the Pólya-Gamma Gibbs sampler. This creates a computational bottleneck because the first stage cannot be parallelized and thus does not scale with the number of computational cores. Johnson et al. (2022) proposed a multi-stage PPRB algorithm for a different class of models in which the transient posterior distribution is approximated using non-Bayesian likelihood maximization in the first stage. A sample from the approximated transient posterior distribution can then be corrected using an approximate M-H ratio in later stages (Johnson et al., 2022).
Applying a similar approach to fit SPP models, we can perform fast maximum likelihood estimation for $\boldsymbol{\beta}$ in the first stage using the glm function in R to fit the logistic regression model described in section 2.2.1. Assuming a flat prior for $\boldsymbol{\beta}$ and standard regularity conditions (i.e., the likelihood is continuously twice-differentiable), we can apply large sample theory to obtain the following approximation:
@@
Johnson et al. (2022) showed that the approximate transient posterior distribution (16) allows for approximate cancellations in the later stage M-H ratio and that parameter estimates obtained using this method are close to those obtained through a conventional single-stage MCMC algorithm. This method can be applied in our multi-stage procedure to obtain an approximate and fast sample from the posterior distribution for $\boldsymbol{\beta}$. Using the Gibbs sampler in (14), the M-H ratio in the second stage again reduces to (15). We refer to this multi-stage method as the GLM approximate (GLM-A) method hereafter.
3.2.1 GLM Exact Approach
If exact inference is desired, we propose an adjustment to the approximate approach developed by Johnson et al. (2022) to make it exact. After performing maximum likelihood estimation and obtaining an approximate transient posterior distribution as in (16), we can obtain an exact M-H correction for the complete likelihood (1) that retains the contribution from the original prior for $\boldsymbol{\beta}$ and the approximate transient posterior proposal distribution. Using the Gibbs sampler for $\zeta$ based on the full-conditional distribution in (14) results in the following M-H ratio:
@@ 3.2.1 GLM Exact Approach
Note that this approach differs from the previously described PPRB strategies (PG, HMC, and GLM-A) because we use the naive (i.e., noninformative) prior as opposed to the transient posterior distribution as the prior distribution for $\boldsymbol{\beta}$. However, this approach still differs from a traditional MCMC algorithm in that it uses a PPRB-style strategy to obtain an efficient proposal distribution and a multi-stage construction to leverage parallel computing resources.
To accelerate the evaluation of the log-likelihood for $\big[\{\mathbf{s}_{i}\},n|{\beta_{0}}^{(*)},\boldsymbol{\beta}^{(*)}\big]$ in the second stage, we compute $\sum_{i=1}^{n}\mathbf{x}^{\prime}(\mathbf{s}_{i})\boldsymbol{\beta}$ in parallel in the intermediate stage in addition to $\Lambda(\mathcal{S})$ using the first stage transient posterior sample. Although the intermediate and second stages involve more calculations compared to the previous multi-stage methods (PG, HMC, and GLM-A), the overall algorithm results in significant speedup for this model compared to a conventional single-stage MCMC algorithm because of the fast first-stage evaluation using maximum likelihood estimation and the ability to parallelize large computations over $k$ in the intermediate stage. We refer to this multi-stage method as the GLM exact (GLM-E) method hereafter.
Both the GLM-A and GLM-E methods result in significant speedup when compared to both the single-stage MCMC algorithm and fully Bayesian multi-stage methods (PG and HMC) due to fast maximum likelihood estimation in the first stage using the glm function in R. Furthermore, the independent samples from the first stage can help prevent sample degeneracy, a common issue that arises in multi-stage MCMC and other Bayesian filtering algorithms (Barreto and Hooten, 2025; Scharf, 2025; Tyalor et al., 2025). Figure 1 provides a flow chart that compares the various first-stage sampling strategies and Table 1 compares the strengths of each approach.
Figure 1: Flowchart of multi-stage algorithm with various first-stage sampling strategies (PG, HMC, GLM-A, GLM-E).
Property PG HMC GLM-E GLM-A
@@
Exact Bayesian inference $\times$
Independent second stage proposals $\times$ $\times$
@@ 3.3.2 Posterior Point Simulation
We can simulate posterior predictive event locations in the unobserved region $\mathcal{S}_{0}$ after obtaining estimates for $\lambda(\mathbf{u})$ $\forall\mathbf{u}:\mathbf{u}\in\mathcal{S}_{0}$ using the Lewis-Shedler method for simulating from an IPP (Lewis and Shedler, 1979). Alternatively, a new point realization for the entire study domain $\mathcal{S}$ can be simulated using posterior estimates for $\lambda(\mathbf{u})$ $\forall\mathbf{u}:\mathbf{u}\in\mathcal{S}$. These results can then be used to characterize posterior predictive derived quantities or functions, such as the Ripley’s K function or pair correlation function.
@@ 4 Simulation Study
We demonstrate the multi-stage MCMC approach using a simulation study in a compact window setting. The data were generated from an IPP with an intensity function that is log-linear in terms of three known parameters across the entire study domain (Figure 2). Only the events simulated within the predefined compact windows were used to fit the model with the windowed complete likelihood (3).
Figure 2: (a) Intensity heat map for simulated data. (b) Realization of simulated points. Points located within the compact windows (red; $n$ = 580) were used to fit the model.
We fit our model to the simulated data using the first-stage sampling strategies described in section 3.2 and a conventional single-stage MCMC algorithm for comparison on a 24-core machine with 3.68 GHz processors and 192 GB of RAM. Figure 3 compares the posterior samples for $\beta_{0}$ and $\boldsymbol{\beta}$ obtained using the single-stage and the multi-stage MCMC algorithms. It is evident that the marginal posterior distributions for $\beta_{0}$ and $\boldsymbol{\beta}$ are almost identical among implementation approaches, barring Monte Carlo error and the approximation error from the logistic regression approximation. Similarly, the joint posterior distributions for each pair of model parameters appear nearly indistinguishable between the single-stage method and the GLM-E method (Figure 4).
@@
Figure 4: Comparison of bivariate joint posterior distributions for each pair of model parameters. Contours represent kernel density estimates of the joint posterior distributions. The GLM-E contours (pink) are overlaid on top of the single-stage contours (gray).
It is worth noting that while the PG, HMC, and GLM-A methods experience some, but often negligible, approximation error from the logistic regression approximation in the first stage, the GLM-E method is unaffected because it does not approximate the conditional likelihood (2) with logistic regression in the M-H correction. Increasing the number of background points $m$ can decrease the logistic regression approximation error, but this incurs significant computational costs for the PG and HMC methods.
Table 2 compares the seconds per effective sample size between different methods for each parameter. The time recorded for the single-stage method does not include a priori parameter tuning, while the remaining multi-stage methods did not require tuning. The Bayesian first-stage methods (PG and HMC) showed an order of magnitude improvement, while the non-Bayesian first-stage methods (GLM-E and GLM-A) showed two orders of magnitude improvement. These results were due to both an improvement in runtime and MCMC effective sample size. All algorithms were fit using $L=640,000$ quadrature points and the results scale with $L$ and the number of computational cores. Figure 5 shows the posterior predictive distribution for $N$ obtained using the posterior samples from the GLM-E multi-stage method. Similar results were obtained for three additional simulations with varying compact window sizes and latent intensity fields (Appendix B).
@@ 5 Application: Harbor Seal Pups in Johns Hopkins Inlet
Previous studies analyzing these data have focused on modeling harbor seal abundance for the entire inlet or using coarse spatial grids. For example, Womble et al. (2020) incorporated the aerial imagery survey data along with other sources of data to obtain estimates and uncertainty quantification for harbor seal abundance. Womble et al. (2021) fit a multivariate conditional autoregressive (MCAR) model to summarize spatially explicit abundance on a 200 m $\times$ 300 m grid ($L=434$ total grid cells). Jansen et al. (2015) used harbor seal data from a similar aerial survey to fit a generalized additive model (GAM) and summarized spatially explicit abundance on a grid of 400 m $\times$ 400 m cells ($L=779$ total grid cells). Previous studies on spatially explicit abundance of harbor seals in Johns Hopkins Inlet have not explored modeling the locations as a spatial point process directly. Modeling these data as an SPP provides unique advantages because it allows for informed simulation of potential seal locations outside the image footprints using an established statistical framework. Furthermore, the multi-stage algorithm makes fitting the model on a fine spatial grid more computationally flexible and feasible, especially with the availability of distributed computing resources.
We fit a Bayesian SPP model to harbor seal pup haul out location data from 21 June 2007 ($n=452$; Figure 7) using the aforementioned multi-stage MCMC approaches. We modeled harbor seal pups because they are the most vulnerable segment of the population and are especially susceptible to disturbance from vessels (Mathews et al., 2016). Furthermore, pups are hauled out on ice and observable with high probability at the time the survey was conducted (i.e., in the month of June and in the early afternoon) based on secondary analysis of wet-dry tag data of harbor seal pups in Disenchantment Bay, Alaska. Our study domain spanned from the Johns Hopkins Glacier terminus to Jaw Point, which is approximately 9km long and 1.5-2km wide (Figure 6a). The image footprints for this survey cover approximately half of the study domain. We assume perfect detection of hauled out pups within the image footprints and no detection outside the image footprints, giving rise to compact observation windows. Additionally, due to the systematic surveying within the study domain, we assume the unobserved areas in the inlet to be ignorable missing data (i.e., the covariate values and relationship between covariates and pup locations do not differ significantly between observed and unobserved areas). Using the pup locations as point events, we fit the SPP model using the windowed complete likelihood (3).
@@
We fit the model using three covariates: ice proportion, bathymetry, and distance to glacier terminus (Appendix C). The ice data were represented as iceberg outlines within the image footprints and summarized as ice proportions to allow for spatial prediction outside the image footprints (Appendix D). Each covariate was summarized on the same raster grid with a spatial resolution of approximately 10 m $\times$ 10 m ($L=304,945$ total grid cells). The grid centers were used as quadrature points throughout the analysis.
5.1 Accounting for Additional Spatial Heterogeneity
From preliminary analysis and model checking, we found that using a log-linear combination of the three covariates to model the IPP intensity was insufficient in capturing the heterogeneity of the observed harbor seal pup data (Appendix E). We first discuss how we modified the model to account for additional spatial heterogeneity for the harbor seal pup model in addition to alternative strategies and their compatibility with the multi-stage algorithm.
5.1.1 Neural Network Basis Expansion
To increase model flexibility and allow for more complex relationships between covariates and intensity, we constructed a basis representation of $\mathbf{X}$, the $p$ $\times$ $L$ design matrix consisting of the covariate values at each grid center. Such practice is common in spatial modeling, with the popularity of fixed-rank (Cressie and Johannesson, 2008) and radial (Buhmann, 2000) basis functions. Of the basis functions typically used to account for spatial heterogeneity (Hefley et al., 2017; Wikle and Zammit-Mangion, 2023), we found that a single-hidden layer feedforward neural network (SLFNN) was able to capture the relationship between the covariates and pup space use. Additionally, we found that the Extreme Learning Machine (ELM) scheme (Huang et al., 2006) was sufficient for constructing the neural network (NN) weights and resulted in efficient computation in the multi-stage algorithm.